OpenStudy (anonymous):

(x)/(x-2) - (-2)/(x+2)=(8)/(4-x^2)

hero (hero):

\[\frac{x}{x - 2} - \frac{-2}{x + 2} = \frac{8}{4 - x^2}\] For convenience express the middle fraction as an addition: \[\frac{x}{x - 2} + \frac{2}{x + 2} = \frac{8}{4 - x^2}\] And make sure the denominators are factored: \[\frac{x}{x - 2} + \frac{2}{x + 2} = \frac{8}{(2 + x)(2 - x)}\]

OpenStudy (anonymous):

I know this but how do i reverse the (2-x) ?

hero (hero):

Okay to reverse that, take out a negative like so: \[\frac{x}{x - 2} + \frac{2}{x + 2} = -\frac{8}{(x+2)(x-2)}\]

OpenStudy (anonymous):

Then I add this on both sides.

hero (hero):

Here's what to do...

hero (hero):

Add the fraction on the right to both sides; subtract 2/(x + 2) from both sides: \[\frac{x}{x - 2} +\frac{8}{(x+2)(x-2)}= -\frac{2}{x + 2} \]

OpenStudy (anonymous):

Why ?

hero (hero):

Actually...

hero (hero):

Instead of doing that...

hero (hero):

Lets reduce x/(x - 2)

hero (hero):

It's too bad you found out that way. It is better to discover such things on your own.

hero (hero):

I'm going to continue posting my solution anyway.

OpenStudy (anonymous):

Thank you.

hero (hero):

You should try to figure it out on your own why no solutions exist.

OpenStudy (anonymous):

The result is \[\frac{ x+2 }{ x-2 }\]

OpenStudy (anonymous):

I think.

OpenStudy (anonymous):

*=0